OpenStudy (anonymous):

i need help on a word problem please this is it you are running a small business and your accountant says you must set aside 22% of total revenue for taxes. you spend another 3600 per month on other expenses. how much total revenue do you need to generate in order to clear a profit of 6000 per month

OpenStudy (anonymous):

R - E = P is a general equation to find profit (revenue minus expenses). You must subtract 22% of your R or .22R from your R. R - .22R - E = P .78R - E = P E is 3600 per month and you want P to be 6000 per month. .78R - 3600 = 6000 .78R = 9600 R = 12307.69 You must generate about 12304.69 in total revenue.